what are the points on a graph if y=2x+2;y=x-3

Answer by ewatrrr(24785)


Hi
the following are linear equations representing lines on a graph
I. Can find points(x,y) on these lines by sustituting values of x
y = 2x+2 |Pt(0,2) and Pt(1,4) are two points on this line
y = x-3 |Pt(0,-3) and Pt(1,-2) are two points on this line
II.Plotting these points, one can the draw lines thru them representing the equation
for example: green is y = 2x+2 and blue is y = x-3